The standard deviation of the market-index portfolio is 45%. Stock A has a beta of 1.20 and a residual standard deviation of 50%.


a. Calculate the total variance for an increase of 0.20 in its beta. (Do not round intermediate calculations. Round your answer to the nearest whole number.)



b. Calculate the total variance for an increase of 9.61% in its residual standard deviation. (Do not round intermediate calculations. Round your answer to the nearest whole number.)
Answer #1

Total variance=(beta*standard deviation of market)^2+(residual standard deviation)^2

1.
=(1.4*45%)^2+(50%)^2=0.6469

2.
=(1.2*45%)^2+(50%+9.61%)^2=0.64693521